OpenStudy (anonymous):

Get a medal :3 What is the value of the expression 3/3^-2 ? A. 27 B. 18 C. 1/2 D. 1/3

OpenStudy (anonymous):

\[\frac{ 1 }{ a^{-1} } = 1\div \frac{ 1 }{ a } \rightarrow 1 \times \frac{ a }{ 1 } = a\]

OpenStudy (anonymous):

We can use that basic idea for your problem now\[\frac{ 3 }{ 3^{-2}} = 3 \div \frac{ 1 }{ 3^2 } = 3 \times 3^2 = 27\]

OpenStudy (anonymous):

@d3Xter Ok, thank you for explaining this to me :3 Here's your medal :)

OpenStudy (anonymous):

You're welcome. :)